OpenStudy (anonymous):

Find the radius of a circle with the given circumference.use 3.14 for pi.C=142.87cm.(Also I'm leaving to go somewhere so it would be nice if you could solve that for me.

OpenStudy (kkutie7):

\[C=2*3.14*r\] \[142.87=2*3.14*r\] \[\frac{142.87}{2*3.14}=r\] you can calculate from here